# Optimal Monotone Depth-Three Circuit Lower Bounds for Majority

### v1

@@ 4.6.3 There exists a 2-set that is disjoint from the other two 2-sets
One of $\mathscr{P}(3,1),\mathscr{P}(3,2)$ and $\mathscr{P}(3,3)$ always holds for every set system in $\Xi$}.
4.6.3.1 $s$ is odd and the other two sets $\{c,d\}$ and $\{e,f\}$ overlap  – $\mathscr{P}(3,1)$
If $\{c,d\}$ and $\{e,f\}$ overlap, the following calculation applies.
$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
1 n1n-1 t1t-1 s2s-2 $2o$ $\frac{5}{9}$
0 1 n2n-2 t1t-1 s1s-1 $2o$ $\frac{4}{9}$
@@ 4.6.3.1 $s$ is odd and the other two sets $\{c,d\}$ and $\{e,f\}$ overlap  – $\mathscr{P}(3,1)$
4.6.3.2 $s$ is odd, $a$ appears in no other set and $\{c,d\}\cap\{e,f\}=\varnothing$  – $\mathscr{P}(3,2)$
$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
1 0 n2n-2 t1t-1 s1s-1 $2d$ $\frac{25}{54}$
0 1 n2n-2 t1t-1 s1s-1 $2d$ $\frac{25}{54}$
Total $\frac{25}{27}$
4.6.3.3 $s$ is odd, $a$ appears in some other 3-set and $\{c,d\}\cap\{e,f\}=\varnothing$  – $\mathscr{P}(3,3)$
$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
1 n1n-1 t1t-1 s2s-2 $2d$ $\frac{5}{9}$
0 1 n2n-2 t1t-1 s1s-1 33 $\frac{7}{18}$
Total $\frac{17}{18}$
We now deal with the case that $s$ is even. Let $\Xi$} be the class of such 3-set systems. We cover $\Xi$} based on the following properties.
$\mathscr{P}(3,1)$:
$s$ is even and either $a$ or $b$ appear in no other set.
$\mathscr{P}(3,2)$:
$s$ is even and either $a$ or $b$ appears in a 3-set with a variable from $\{c,d,e,f\}$.
$\mathscr{P}(3,3)$:
$s$ is even, $|\{c,d,e,f\}|=3$ and there exists set $\{v_{ab},x,y\}$ where $v_{ab}\in\{a,b\}$ and $\{x,y\}\cap\{c,d,e,f\}=\varnothing$.
$\mathscr{P}(3,4)$:
$s$ is even, $c,d,e,f$ are distinct, there exist $v_{ab}\in\{a,b\}$ and $v_{cdef}\in\{c,d,e,f\}$ such that there exist sets $\{v_{ab},i,j\}$ and $\{v_{cdef},x,y\}$ where $\{x,y\}\neq\{i,j\}$ and none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
$\mathscr{P}(3,5)$:
$s$ is even, $c,d,e,f$ are distinct, there exists $v_{ab}\in\{a,b\}$ such that there exist sets $\{v_{ab},i,j\}$, $\{v_{ab},x,y\}$ where none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
$\mathscr{P}(3,6)$:
$s$ is even, $c,d,e,f$ are distinct, and there exist $x,y$ such that for all $v_{abcdef}\in\{a,b,c,d,e,f\}$, the set $\{v_{abcdef},x,y\}$ exists.
Lemma 4.12.
For $i\in[6]$, one of $\mathscr{P}(3,i)$ always holds for every set system in $\Xi$}.
Proof.
Let $F\in{\Xi}$ be arbitrary and assume none of $\mathscr{P}(3,i)$ holds for any choice of isolated 2-set in the three 2-set system.
Since $\mathscr{P}(3,1)$ and $\mathscr{P}(3,2)$ don’t hold, $a$ and $b$ appear in at least one other clause, and all such clauses don’t contain any of $c,d,e,f$. Notice that amongst $c,d,e,f$, by non-redundancy, either there exist $3$ distinct variables or $4$. If there were $3$, then we see that $\mathscr{P}(3,3)$ must have been satisfied. Hence, it must be the case that $c,d,e,f$ are all distinct.
This means that the three 3-sets in $F$ must be disjoint. Therefore by symmetry, instead of letting $\{a,b\}$ be the isolated 2-set, we can relabel variables and let any of $\{c,d\}$ or $\{e,f\}$ be the isolated 2-set and check for these properties. Hence, we can conclude that none of $a,b,c,d,e,f$ appears in any other set with each other (this is ruled out already for say $a,b$ by non-redundancy and remaining by negation of $\mathscr{P}(3,2)$ and symmetry). Also, we must have that each of $a,b,c,d,e,f$ must appear in at least one more 3-set. Since $\mathscr{P}(3,5)$ does not hold, it must be that both $a,b$ appear in exactly one other clause and by symmetry that each of $c,d,e,f$ appear in exactly one other clause. Also, since $\mathscr{P}(3,4)$ does not hold, we conclude that if the unique other 3-set where $a$ appears is say $\{a,x,y\}$ then it must be the case that there exists 3-set $\{c,x,y\}$. Symmetrically applying this argument to all pairs, we infer that case $\mathscr{P}(3,6)$ must hold for $F$, which is a contradiction. ∎
4.6.3.4 $s$ is even and either $a$ or $b$ appear in no other set  – $\mathscr{P}(3,1)$
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
1 0 n2n-2 t1t-1 s1s-1 22 $\frac{10}{21}$
0 1 n2n-2 t1t-1 s1s-1 22 $\frac{10}{21}$
Total $\frac{20}{21}$
4.6.3.5 $s$ is even and either $a$ or $b$ appears in a 3-set with a variable from $\{c,d,e,f\}$  – $\mathscr{P}(3,2)$
Without loss of generality let this be $a$ and $c$. Let this set be $\{a,c,x\}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$
@@ 4.6.3.4 $s$ is even and either $a$ or $b$ appear in no other set  – $\mathscr{P}(3,1)$
4.6.3.6 $s$ is even, $|\{c,d,e,f\}|=3$ and there exists set $\{v_{ab},x,y\}$ where $v_{ab}\in\{a,b\}$ and $\{x,y\}\cap\{c,d,e,f\}=\varnothing$  – $\mathscr{P}(3,3)$
@@
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
@@
4.6.3.7 $s$ is even, $c,d,e,f$ are distinct, there exist $v_{ab}\in\{a,b\}$ and $v_{cdef}\in\{c,d,e,f\}$ such that there exist sets $\{v_{ab},i,j\}$ and $\{v_{cdef},x,y\}$ where $\{x,y\}\neq\{i,j\}$ and none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$  – $\mathscr{P}(3,4)$
@@
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$
@@
4.6.3.8 $s$ is even, $c,d,e,f$ are distinct, there exists $v_{ab}\in\{a,b\}$ such that there exist sets $\{v_{ab},i,j\}$, $\{v_{ab},x,y\}$ where none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$  – $\mathscr{P}(3,5)$
@@
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
@@
4.6.3.9 $s$ is even, $c,d,e,f$ are distinct, and there exist $x,y$ such that for all $v_{abcdef}\in\{a,b,c,d,e,f\}$, the set $\{v_{abcdef},x,y\}$ exists  – $\mathscr{P}(3,6)$
We directly compute:
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$
0 1 n2n-2 t1t-1 s1s-1 33 $\frac{3}{7}$
1 0 n2n-2 t1t-1 s1s-1 33 $\frac{3}{7}$
1 1 $x=y=0,c=d=e=f=1$ n8n-8 t6t-6 s10s-10 0 $\frac{1}{56}$
Total $\frac{7}{8}$
@@ 4.7 Proof of the Upper Bounds for $\Phi_{4}(s,t)$
Let $F=F_{1}F_{2}$ for some monotone $3$-CNF $F_{2}$, where $F_{1}$ is composed of four distinct 2-clauses and one of them is $(ab)$. Let $\Xi$} be the class of such 3-set systems. We cover $\Xi$} based on the following properties. (Note we only care about odd values of $s$, as even values are handled by the three 2-clause case.)
@@
Lemma 4.13.
Some $\mathscr{P}(4,j)$ must hold in a formula with at least have four two-clauses.
Proof.
Proceed by contradiction, assuming $\neg\bigwedge_{j=1}^{3}\mathscr{P}(4,j)$. In this case, by $\neg\mathscr{P}(4,1)$ some two-clause intersects another. This means some variable is in more than one 2-clause. By $\neg\mathscr{P}(4,2)$, it must be in at least three 2-clauses, implying $\mathscr{P}(4,3)$. ∎
4.7.0.1 No two two-clauses intersect  – $\mathscr{P}(4,1)$
In this case, no three of the 2-clauses form a star, as they don’t overlap at all. We can therefore use the standard type 3 bound.
$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
@@ 4.7.0.1 No two two-clauses intersect  – $\mathscr{P}(4,1)$
4.7.0.2 Some variable is in exactly two 2-clauses  – $\mathscr{P}(4,2)$
Without loss of generality, assume $a$ is in both $ab$ and $ac$. Note $b$ and $c$ can be in as many 2-clauses as desired.
$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$
@@ 4.7.0.2 Some variable is in exactly two 2-clauses  – $\mathscr{P}(4,2)$
4.7.0.3 Some variable is in at least three 2-clauses  – $\mathscr{P}(4,3)$
Without loss of generality, assume $a$ is in $ab$, acac, and $ad$. It may be in as many other 2-clauses as desired, as may $b$, $c$, and $d$.
$a$ $b$ Consequence $n$ $t$ $s$ Type Odd $s$

### v2

@@ 0 $b=1$ n2n-2 t1t-1 s1s-1 33 $\frac{7}{18}$
For $i\in[6]$, one of $\mathscr{P}(3\text{-}{ie},i)$ always holds for every monotone $3$-CNF in ${\Xi}_{3\text{-}i}(s,t)$ when $s$ is even.
Proof.
For even $s$, let $F\in{\Xi}_{3\text{-}i}(s,t)$ be arbitrary and assume none of $\mathscr{P}(3\text{-}{ie},i)$ holds for any choice of isolated $2$-clause in the three $2$-clause set.
Since $\mathscr{P}(3\text{-}{ie},1)$ and $\mathscr{P}(3\text{-}{ie},2)$ don’t hold, $a$ and $b$ appear in at least one other clause, and all such clauses don’t contain any of $c,d,e,f$. Notice that among $c,d,e,f$, by non-redundancy, either there exist $3$ distinct variables or $4$. If there were $3$, then we see that $\mathscr{P}(3\text{-}{ie},4)$ must have been satisfied. Hence, it must be the case that $c,d,e,f$ are all distinct.
This means that the three 2-sets in $F$ must be disjoint. Therefore by symmetry, instead of letting $\{a,b\}$ be the isolated 2-set, we can relabel variables and let any of $\{c,d\}$ or $\{e,f\}$ be the isolated 2-set and check for these properties. Hence, we can conclude that none of $a,b,c,d,e,f$ appears in any other set with each other (this is ruled out already for say $a,b$ by non-redundancy and remaining by negation of $\mathscr{P}(3\text{-}{ie},2)$ and symmetry). Also, we must have that each of $a,b,c,d,e,f$ must appear in at least one more 3-set. Since $\mathscr{P}(3\text{-}{ie},6)$ does not hold, it must be that both $a,b$ appear in exactly one other clause and by symmetry that each of $c,d,e,f$ appear in exactly one other clause. Also, since $\mathscr{P}(3\text{-}{ie},5)$ does not hold, we conclude that if the unique other 3-set where $a$ appears is say $\{a,x,y\}$ then it must be the case that there exists 3-set $\{c,x,y\}$. Symmetrically applying this argument to all pairs, we infer that case $\mathscr{P}(3\text{-}{ie},7)$ must hold for $F$, which is a contradiction. ∎
6.6.3.4 $\mathscr{P}(3\text{-}{ie},1)$$s$ is even and either $a$ or $b$ appear in no other clause.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
1 $b=0$ n2n-2 t1t-1 s1s-1 22 $\frac{10}{21}$
0 $b=1$ n2n-2 t1t-1 s1s-1 22 $\frac{10}{21}$
Total $\frac{20}{21}$
6.6.3.5 $\mathscr{P}(3\text{-}{ie},2)$$s$ is even, either $a$ or $b$ appears in a 3-clause with a variable from $\{c,d,e,f\}$, and that variable is in at most one of $cd$ and $ef$.
Without loss of generality, let this be $a$ and $c$. Let this set be $\{a,c,x\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
1 n1n-1 t1t-1 s2s-2 22 $\frac{25}{42}$
0 1 $b=1$ n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$
0 0 $b=d=x=1$ n5n-5 t3t-3 s4s-4 11 $\frac{5}{42}$
@@ 0 $b=1$ n2n-2 t1t-1 s1s-1 22 $\frac{10}{21}$
6.6.3.6 $\mathscr{P}(3\text{-}{ie},3)$$s$ is even, either $a$ or $b$ appears in a 3-clause with a variable from $\{c,d,e,f\}$, and that variable is in both of $cd$ and $ef$.
Without loss of generality let this be $a$ and $c$, where $c=f$ and the additional two-clauses are $cd$, cece. Let this set be $\{a,c,x\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
@@
6.6.3.7 $\mathscr{P}(3\text{-}{ie},4)$$s$ is even, $|\{c,d,e,f\}|=3$ and there exists clause $\{v_{ab},x,y\}$ where $v_{ab}\in\{a,b\}$ and $\{x,y\}\cap\{c,d,e,f\}=\varnothing$.
@@
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
@@ 0 $b=1$ n2n-2 t1t-1 s1s-1 33 $\frac{3}{7}$
6.6.3.8 $\mathscr{P}(3\text{-}{ie},5)$$s$ is even, $c,d,e,f$ are distinct, there exist $v_{ab}\in\{a,b\}$ and $v_{cdef}\in\{c,d,e,f\}$ such that there exist clauses $\{v_{ab},i,j\}$ and $\{v_{cdef},x,y\}$ where $\{x,y\}\neq\{i,j\}$ and none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
@@
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
@@
6.6.3.9 $\mathscr{P}(3\text{-}{ie},6)$$s$ is even, $c,d,e,f$ are distinct, there exists $v_{ab}\in\{a,b\}$ such that there exist clauses $\{v_{ab},i,j\}$, $\{v_{ab},x,y\}$ where none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
@@
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$
@@ 6.7 Proof of the Upper Bounds for $\Phi_{4}(s,t)$
Let $\mathbb{F}=(V,F_{1}\cap F_{2})\in{\Xi}_{4}(s,t)$ be a monotone $3$-CNF over $|V|$ variables with threshold $t$ and deficit $s$ where $F_{1}$ contains exactly four distinct 2-clauses and one of them is $\{a,b\}$. We cover $\Xi$} based on the following properties. Note we only care about odd values of $s$, as even values are handled by the three 2-clause case.
@@
6.7.1 $\mathscr{P}(4,1)$ – No two $2$-clauses intersect
Assume $\mathbb{F}\in{\Xi}_{4}(s,t)$ satisfies $\mathscr{P}(4,1)$. In this case, no three of the 2-clauses form a star, as they don’t overlap at all. We can therefore use the standard type 3 bound.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
@@ 0 $b=1$ n2n-2 t1t-1 s1s-1 33 $\frac{7}{17}$
6.7.2 $\mathscr{P}(4,2)$ – Some variable is in exactly two 2-clauses
Assume $\mathbb{F}\in{\Xi}_{4}(s,t)$ satisfies $\mathscr{P}(4,3)$. Without loss of generality, assume $a$ is in both $ab$ and $ac$. Note $b$ and $c$ can be in as many 2-clauses as desired.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
@@ 6.7.2 $\mathscr{P}(4,2)$ – Some variable is in exactly two 2-clauses
6.7.3 $\mathscr{P}(4,3)$ – Some variable is in at least three 2-clauses
Assume $\mathbb{F}\in{\Xi}_{4}(s,t)$ satisfies $\mathscr{P}(4,3)$. Without loss of generality, assume $a$ is in $\{a,b\}$, $\{a,c\}$, and $\{a,d\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
@@ Appendix A Optimality of Building Blocks
Notation A.1.
For a monotone CNF $\mathbb{F}=(V,F)$ and for $S\subset V$, we let $\mathbb{F}^{S}=(V-S,\mathbb{F}^{S})$ denote the CNF obtained by deleting the sets in $F$ that contain a variable in $S$, that is, $F^{S}=\{C\in F\mid C\cap S=\varnothing\}$. $\mathbb{F}^{S}$ corresponds to the CNF obtained by setting the variables in $S$ to 1. We use $\mathbb{F}^{\neg S}=(V-S,F^{\neg S})$ to denote the CNF obtained by deleting all occurrences of variables in $S$ from the clauses in $F$, that is, $F^{\neg S}=\{C-S\mid C\in F\}$. $\mathbb{F}^{\neg S}$ refers to the CNF obtained by setting the variables in $S$ to 0. For a variable $x$, let $\mathbb{F}^{x}=\mathbb{F}^{\{x\}}$ and $\mathbb{F}^{\neg x}=\mathbb{H}^{\neg\{x\}}$. For disjoint sets of variables $S$ and $T$, we use $\mathbb{F}^{S^{\prime},T^{\prime}}$ to denote be $S$ or $\neg S$ and $T^{\prime}$ can be $T$ or $\neg T$.